OpenStudy (anonymous):

log(base 16) x + log(base 4) x + log(base 2) x = 7. Solve for X

OpenStudy (anonymous):

\[\log_{16} x + \log_{4} x + \log_{2} x = 7\]

OpenStudy (jkristia):

This is how I solved it, I believe it is correct, but I'm sure there is an easier way. First convert to same base \[\log _{16}x + \log_{4}x+\log_{2}x = 7\]\[\log_{2}x^{\frac{1}{4}}+\log_{2}x^{\frac{1}{2}} + \log_{2}x = 7\] The combine into one log \[\log_{2}(x^{\frac{1}{4}}+x^{\frac{1}{2}} + x ^1)= 7 => \log_{2}(x^{\frac{7}{4}})= 7\] This means \[2^7 = x^{\frac{7}{4}}\] Raise both sides by 4/7 (to get x^1) \[2^{7{\frac{4}{7}}} = x^1=> 2^4 = x => x = 16\]

OpenStudy (jkristia):

oops, there is a typo. When combining the log the x's should be multiplied, not added.